OpenStudy (anonymous):

70% of what number is 224?

OpenStudy (xapproachesinfinity):

70%x=224 solve for x

OpenStudy (xapproachesinfinity):

to make it easier here a tip replace of with multiplication what number meaning the number you looking for name it some letter "is" replace it with equal sign and then solve for the letter whatever it is you choose

OpenStudy (xapproachesinfinity):

remember 70% is written 70/100

OpenStudy (anonymous):

This should be very easy. Multiplying a number by 1 means getting its 100 percent. Multiplying a number by 2 means getting its 200 percent. Also, by multiplying a number 1.5 means getting its 150 percent. Therefore, by multiplying that number with 0.70 means getting its 70 percent. I hope you got it :)

OpenStudy (xapproachesinfinity):

here we are not multiplying we are dividing since we are looking for that number^_^

OpenStudy (anonymous):

^ you talkin to me?

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (anonymous):

70/100 = 0.7 (and that's the simplified form) Hence, dividing has already been done. And its much easier since our denominator is 10. Therefore, by multiplying that number by 0.70 which is also the same as 70/100 will get the value. Anyway, there's no difference with the final answers.

OpenStudy (xapproachesinfinity):

okay can you do it that way and tell me the answer?

OpenStudy (xapproachesinfinity):

i wasn't talking about what 70/100

OpenStudy (anonymous):

thanks

OpenStudy (xapproachesinfinity):

wrong!

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

i misread the question

OpenStudy (xapproachesinfinity):

you see you didn't pay attention to im saying

OpenStudy (xapproachesinfinity):

that is what i meant by are dividing nor multiplying

OpenStudy (anonymous):

well that's still easy 0.7x = 224

OpenStudy (xapproachesinfinity):

why*

OpenStudy (anonymous):

hence, x = 224/0.7

OpenStudy (xapproachesinfinity):

yeah that's what i meant we are dividing if we are looking to the number that give 224 as 70 percent

OpenStudy (anonymous):

Okay. I just misread the que

OpenStudy (xapproachesinfinity):

No worries^_^ i think our friend got the idea here

OpenStudy (xapproachesinfinity):

i terrible miss spelled everything lol

OpenStudy (xapproachesinfinity):

terribly*

OpenStudy (midhun.madhu1987):